OpenStudy (anonymous):

find the slope of the line: (1/3)x - (-1/5)y=1

OpenStudy (anonymous):

That equation can be re-written as: (1/3)x + (1/5)y = 1 -> (1/5)y = (-1/3)x + 1 y = (-5/3)x + 5 That is in the form y = mx + b where (-5/3) is "m" or the slope you are looking for.

OpenStudy (anonymous):

i tried and i was way off, thank you!

OpenStudy (anonymous):

Thx again. np. We are always here to help. Someone always shows up! And you're welcome!